A railway track (made of iron) is laid in winter when the average temperature is 18°C. The track consists of sections of 12"0 m placed one after the other. How much gap should be left between two such sections so that there is no compression during summer when the maximum temperature goes to 48°C ? Coefficient of linear expansion of iron = 11 x 10 6/°C.

Grade:11

1 Answers

Gaurav
6 years ago
As track will expand in summer so as to avoid compression between the track
Length of gap required=2l0\alpha\Deltat=2*12*11*10-6(48-18)
                                           =7.92*10-4 m